### v1

# Non-existence of Ulrich modules over Cohen-Macaulay local rings

Abstract.
Over a Cohen-Macaulay local ring, the minimal number of generators of a maximal Cohen-Macaulay module is bounded above by its multiplicity. In 1984 Ulrich asked whether there always exist modules for which equality holds; such modules are known nowadays as Ulrich modules. We answer this question in the negative by constructing families of two dimensional Cohen-Macaulay local rings that have no Ulrich modules. Some of these examples are Gorenstein normal domains; others are even complete intersection domains, though not normal.
Key words and phrases:
Cohen-Macaulay ring, section ring, Ulrich module, Ulrich sheaf
2020 Mathematics Subject Classification:
13C13 (primary); 13H10, 13C14, 14F06 (secondary)
1. Introduction
An Ulrich module over a commutative noetherian local ring $R$ is a nonzero maximum Cohen-Macaulay (MCM) $R$-module $U$ with the property that its minimal number of generators, $\nu_{R}(M)$, is equal to is Hilbert-Samuel multiplicity, $e_{R}(U)$. In [2] such modules are referred to as “maximally generated” MCM modules because the inequality $\nu_{R}(M)\leq e_{R}(M)$ holds for any MCM $R$-module $M$.
The mere existence of an Ulrich $R$-module has striking implications. For instance, it implies that Lech’s conjecture holds for $R$: For any flat local map $R\to S$ of local rings, one has $e(R)\leq e(S)$; see, for instance, [8, pg. 69], or [15, Theorem 2.11]. Moreover, if $R$ admits an Ulrich module with the additional property that its class in the Grothendieck group coincides with the class of a free module, then for any $R$-module $L$ of finite length and finite projective dimension, one has $\operatorname{length}(L)\geq e(R)$; see [12] for more general results.
The existence of Ulrich modules is known in the following cases.
  1. $\bullet$
    When $\dim R\leq 1$. For $\dim(R)=0$, the residue field $R/\mathfrak{m}$ is Ulrich and for $\dim R=1$, the ideal $\mathfrak{m}^{e(R)-1}$ is Ulrich as an $R$-module [2, 2.1].
  2. $\bullet$
    Hypersurfaces and, more generally, local rings $R$ such that $\operatorname{gr}_{\mathfrak{m}}(R)$ is a complete intersection [10].
  3. $\bullet$
    When $\dim R\leq 2$ and $R$ is a standard graded ring over a field; the Cohen-Macaulay case is in [2] and the general case follows from [4].
  4. $\bullet$
    Complete local rings $R$ with $\dim R\leq 2$ and $\operatorname{Proj}(\operatorname{gr}_{\mathfrak{m}}(R))$ geometrically reduced. This result will appear in forthcoming work by the first three authors.
See E. Coskun’s article [3] for a nice survey on Ulrich modules and Ulrich sheaves.
Ulrich asked [20, §3] whether every Cohen-Macaulay local ring admits an Ulrich [^1] He did not refer to them as “Ulrich modules”, of course. module. Yhee [21] constructed local domains of dimension two that have no Ulrich modules. However, these examples are not Cohen-Macaulay.
In this work, we answer Ulrich’s original question in the negative:
Theorem 1.1.
There exist Gorenstein normal domains, and complete intersection local domains, of dimension two that have no Ulrich modules.
This result packages Theorems 4.1 and 5.1.
Our technique for producing such examples involves a relationship between Ulrich modules on $R$ and Ulrich sheaves on the projective scheme $E\coloneqq\operatorname{Proj}(\operatorname{gr}_{\mathfrak{m}}(R))$. Specifically, if $U$ is an Ulrich module, then the coherent sheaf $\mathcal{F}$ on $E$ associated to the graded module $\operatorname{gr}_{\mathfrak{m}}U$ is an Ulrich sheaf. When $R$ has dimension two, $E$ is a curve and every curve admits an Ulrich sheaf, and thus at first glance this approach seems unpromising. However, $\mathcal{F}$ extends to the blowup of $\operatorname{Spec}R$ at $\mathfrak{m}$. For the rings occurring in Theorem 1.1, we show no such “extendable” Ulrich sheaves exist on $E$. See Section 2 for details.
Our examples can be obtained by starting with an ample line bundle $\mathcal{L}$ on a projective curve $C$ such that $\Gamma(C,\mathcal{L})=0$ and the canonical pairing
$\Gamma(C,\mathcal{L}^{2})\otimes_{k}\Gamma(C,\mathcal{L}^{j})\longrightarrow\Gamma(C,\mathcal{L}^{j+2})$
is a surjection for all $j\geq 2$. In this situation, the local ring $R$ obtained by localizing (or completing) the section ring $S\coloneqq\bigoplus_{j\geq 0}\Gamma(C,\mathcal{L}^{j})$ at its homogeneous maximal ideal, cannot have any Ulrich modules; see Corollary 3.2.
The simplest situation in which this occurs is when $C$ is a smooth quartic plane curve and $\mathcal{L}$ is a non-effective theta characteristic. There are $36$ choices for such an $\mathcal{L}$ (assuming $k$ is algebraically closed of characteristic not $2$), and each choice leads to a complete Gorenstein normal domain of dimension two without any Ulrich modules. The complete intersection example in Theorem 1.1 arises from the singular projective plane curve $C$ cut out by the polynomial $(y^{3}+x^{2}z)^{2}-x^{3}z^{3}$ and a line bundle $\mathcal{L}$ satisfying $\mathcal{L}^{2}\cong\mathcal{O}_{C}(1)$.
2. Ulrich modules and Ulrich sheaves
Let $(R,\mathfrak{m},k)$ be a noetherian local ring with maximal ideal $\mathfrak{m}$ and residue field $k$. Let $M$ be a nonzero finitely generated $R$-module. The Krull dimension and depth of $M$ are denoted $\dim_{R}M$ and $0pt_{R}M$, respectively. The $R$-module $M$ is maximal Cohen-Macaulay if it is finitely generated and $0pt_{R}M=\dim R$. We write $\nu_{R}(M)$ for the minimal number of generators of $M$, and $e_{R}(M)$ for its Hilbert-Samuel multiplicity with respect to the ideal $\mathfrak{m}$.
Ulrich modules
Any maximal Cohen-Macaulay $R$-module $M$ satisfies $\nu_{R}(M)\leq e_{R}(M)$; see [20, Section 3]. When $e_{R}(M)=\nu_{R}(M)$ and $M$ is nonzero it is said to be an Ulrich module.
Assume $k$ is infinite and let $J\subseteq\mathfrak{m}$ be a minimal reduction of $\mathfrak{m}$. For every maximal Cohen-Macaulay $R$-module $M$, we have
$e(\mathfrak{m},M)=e(J,M)=\ell_{R}(M/JM)\geq\ell(M/\mathfrak{m}M)\,.$
Thus a nonzero maximal Cohen-Macaulay module $M$ is Ulrich if and only if $JM=\mathfrak{m}M$ for one (or equivalently, every) minimal reduction $J$ of $\mathfrak{m}$.
Ulrich sheaves
Let $X$ be a projective scheme over a field and let $\mathcal{O}_{X}(1)$ be a very ample line bundle on $X$. A coherent sheaf $\mathcal{F}$ is called an Ulrich sheaf for $(X,\mathcal{O}_{X}(1))$ if it is nonzero and one of the following equivalent conditions holds:
  1. (1)
    $\operatorname{H}^{i}(X,\mathcal{F}(-j))=0$ for all $1\leq j\leq\dim(X)$, and all $i$.
  2. (2)
    $\operatorname{H}^{i}(X,\mathcal{F}(j))=0$ except when $i=0$, $j\geq 0$ or $i=\dim(X)$, $j\leq-\dim(X)-1$.
Suppose $\pi\colon X\to\mathbb{P}^{n}_{k}$ is a finite map such that $\pi^{*}\mathcal{O}_{\mathbb{P}^{n}_{k}}(1)=\mathcal{O}_{X}(1)$; such a map exists when $k$ is infinite. Then a coherent sheaf $\mathcal{F}$ on $X$ is Ulrich if and only if $\pi_{*}\mathcal{F}\cong\mathcal{O}_{\mathbb{P}^{n}_{k}}^{r}$ for some integer $r\geq 1$; see Eisenbud and Schreyer [4].
The definition of an Ulrich sheaf depends on the very ample line bundle $\mathcal{O}_{X}(1)$. We suppress $\mathcal{O}_{X}(1)$ when it is clear from the context. For instance, when $R$ is a standard graded noetherian ring over a field and $X=\operatorname{Proj}(R)$, we take $\mathcal{O}_{X}(1)$ to be the coherent sheaf associated to $R(1)$.
Remark 2.1.
If $\dim(X)=1$, then $\mathcal{F}$ is an Ulrich sheaf if and only if $\mathcal{F}(-1)$ has no cohomology and $\mathcal{F}\neq 0$. Thus the existence of an Ulrich sheaf for $(X,\mathcal{O}_{X}(1))$ is equivalent to the existence of a nonzero coherent sheaf on $X$ with no cohomology, and the latter property does not depend on the choice of $\mathcal{O}_{X}(1)$.
Rees algebras and blowups
Let $R$ be a local ring and $M$ a finitely generated $R$-module. Consider the graded $R$-module
$\mathscr{R}_{\mathfrak{m}}(M)\coloneqq\bigoplus_{n\geqslant 0}(\mathfrak{m}^{n}M)t^{n}\,.$
Then $\mathscr{R}_{\mathfrak{m}}(R)$ is a graded $R$-algebra, called the Rees ring of $R$ with respect $\mathfrak{m}$, and $\mathscr{R}_{\mathfrak{m}}(M)$ is a finitely generated graded $\mathscr{R}_{\mathfrak{m}}(R)$-module. The canonical map $R\to\mathscr{R}_{\mathfrak{m}}(R)$ induces the morphism
$p\colon\operatorname{Bl}_{\mathfrak{m}}(R)\coloneqq\operatorname{Proj}(\mathscr{R}_{\mathfrak{m}}(R))\longrightarrow\operatorname{Spec}R\,.$
This is the blowup of $\operatorname{Spec}R$ at $\{\mathfrak{m}\}$. The strict transform of an $R$-module $M$ is the coherent sheaf on $\operatorname{Bl}_{\mathfrak{m}}(R)$ associated to the $\mathscr{R}_{\mathfrak{m}}(R)$-module $\mathscr{R}_{\mathfrak{m}}(M)$.
The associated graded module of $M$ is
$\operatorname{gr}_{\mathfrak{m}}(M)=k\otimes_{R}\mathscr{R}_{\mathfrak{m}}(M)=\bigoplus_{n\geqslant 0}(\mathfrak{m}^{n}M/\mathfrak{m}^{n+1}M)t^{n}\,.$
Thus $\operatorname{gr}_{\mathfrak{m}}(R)$ is a standard graded $k$-algebra and the $\operatorname{gr}_{\mathfrak{m}}(R)$-module $\operatorname{gr}_{\mathfrak{m}}(M)$ is finitely generated over $\operatorname{gr}_{\mathfrak{m}}(R)$ by ${\operatorname{gr}_{\mathfrak{m}}(M)}_{0}$.
Set $E=\operatorname{Proj}(\operatorname{gr}_{\mathfrak{m}}(R))$; this is the exceptional fiber of the blowup $p$.
[…]
The strict transform of $M$ can also be described as the quotient $p^{*}M/\Gamma_{E}(p^{*}M)$.
Lemma 2.2.
Let $U$ be an Ulrich $R$-module and $\mathcal{F}$ be the strict transform of $U$. Then $\mathbf{L}{j}^{*}\mathcal{F}\cong j^{*}\mathcal{F}$ in $\operatorname{D}^{b}(\operatorname{coh}E)$ and $j^{*}\mathcal{F}$ is an Ulrich sheaf on $E$.
Proof.
Since $U$ is an Ulrich $R$-module, the graded $\operatorname{gr}_{\mathfrak{m}}(R)$-module $\operatorname{gr}_{\mathfrak{m}}(U)$ is Ulrich, in that it is maximal Cohen-Macaulay, generated in degree 00, and its multiplicity is equal to the number of generators. This is proved in [2, Corollary 1.6] under additional assumptions on $R,U$. Here is an argument in the general case.
The multiplicity and minimal number of generators of $\operatorname{gr}_{\mathfrak{m}}(U)$ as a module over $\operatorname{gr}_{\mathfrak{m}}(R)$ coincide with the corresponding invariants of the $R$-module $U$. Thus the crucial point is that $\operatorname{gr}_{\mathfrak{m}}(U)$ is maximal Cohen-Macaulay. We can assume the residue field of $R$ is infinite, and hence that there exists an $U$-regular sequence, say $\boldsymbol{x}\coloneqq x_{1},\dots,x_{d}$, such that $\boldsymbol{x}U=\mathfrak{m}U$. Since $\boldsymbol{x}$ is $U$-regular, one has an isomorphism of $\operatorname{gr}_{(\boldsymbol{x})}(R)$-modules
$(U/\boldsymbol{x}U)[X_{1},\dots,X_{d}]\xrightarrow{\ \cong\ }\operatorname{gr}_{(\boldsymbol{x})}(U)\quad\text{with $X_{i}\mapsto\overline{x}_{i}$.}$
Here $\overline{x}_{i}$ denotes the residue class of $x_{i}$ in $(\boldsymbol{x})/(\boldsymbol{x})^{2}$. Therefore $\operatorname{gr}_{(\boldsymbol{x})}(U)$ is a maximal Cohen-Macaulay module over $\operatorname{gr}_{(\boldsymbol{x})}(R)$. It remains to observe that this module is isomorphic to $\operatorname{gr}_{\mathfrak{m}}(U)$, since $\boldsymbol{x}U=\mathfrak{m}U$.
It follows that $j^{*}\mathcal{F}$, which is the coherent sheaf associated to $\operatorname{gr}_{\mathfrak{m}}(U)$, is an Ulrich sheaf; see [4, Proposition 2.1]. The coherent sheaf $\mathcal{F}(1)$ is the sheaf associated to $\mathscr{R}_{\mathfrak{m}}(\mathfrak{m}U)$ and the map $\mathcal{F}(1)\to\mathcal{F}$ induced by $t\colon\mathcal{O}_{B}(1)\to\mathcal{O}_{B}$, where $B=\operatorname{Bl}_{\mathfrak{m}}(R)$, arises from the canonical inclusion
$\mathscr{R}_{\mathfrak{m}}(\mathfrak{m}U)=\bigoplus_{i\geqslant 0}\mathfrak{m}^{i+1}U\subseteq\bigoplus_{i}\mathfrak{m}^{i}U=\mathscr{R}_{\mathfrak{m}}(U)\,.$
In particular, the map $\mathcal{F}(1)\to\mathcal{F}$ is injective. This proves that $\mathbf{L}{j}^{*}\mathcal{F}\cong j^{*}\mathcal{F}$. ∎
In the context of the Lemma 2.2, one has also $\mathbf{R}{\pi}_{*}\mathcal{F}\cong U$. Moreover, the result has a converse, and this can be used to construct Ulrich modules over certain two dimensional rings. These results will appear in forthcoming work by the first three authors (we will not use them in the sequel though).
3. Criteria for non-existence
In this section we identify obstructions to the existence of an Ulrich module over a local ring in terms of the geometry of the blowup at its maximal ideal.
Theorem 3.1.
Let $R$ be a noetherian ring and $\mathfrak{m}\subset R$ a maximal ideal with $\dim R_{\mathfrak{m}}\geq 2$. Let $\operatorname{Bl}_{\mathfrak{m}}(R)\to\operatorname{Spec}R$ be the blowup at $\{\mathfrak{m}\}$ and $E$ the exceptional fiber of the blowup. If there exists a locally principal ideal sheaf ${\mathcal{I}}$ on $\operatorname{Bl}_{\mathfrak{m}}(R)$ and an integer $n\geq 2$ such that ${\mathcal{I}}^{n}$ defines $E$ in $\operatorname{Bl}_{\mathfrak{m}}(R)$, then neither the local ring $R_{\mathfrak{m}}$ nor its $\mathfrak{m}$-adic completion has any Ulrich modules.
Proof.
Let $\widehat{R}_{\mathfrak{m}}$ denote the $\mathfrak{m}$-adic completion of $R_{\mathfrak{m}}$. Since the maps $R\to R_{\mathfrak{m}}$ and $R\to\widehat{R}_{\mathfrak{m}}$ are flat, it is easy to verify that the hypotheses carry over to both $R_{\mathfrak{m}}$ and $\widehat{R}_{\mathfrak{m}}$. Thus we can assume in the rest of the argument that $R$ is local, with maximal ideal $\mathfrak{m}$, and that $\dim R\geq 2$.
In what follows we write $B$ instead of $\operatorname{Bl}_{\mathfrak{m}}(R)$. Observe that ${\mathcal{I}}^{n}\cong\mathcal{O}_{B}(1)$. Let $W$ be the subscheme of $B$ cut out by $\mathcal{I}$.
We first consider the case when $\dim R=2$; equivalently, $\dim W=1$.
Assume to the contrary that there exists an Ulrich $R$-module $U$. Let $\mathcal{F}$ be the strict transform of $U$ and $\mathcal{U}$ the sheaf on $E$ associated to $\operatorname{gr}_{\mathfrak{m}}(U)$. We may identify $\mathcal{U}$ with $\mathcal{F}/{\mathcal{I}}^{n}\mathcal{F}$. Setting ${\mathcal{U}}_{j}$ to be the coherent sheaf on $E$ given by ${\mathcal{I}}^{j}\mathcal{F}/{\mathcal{I}}^{n}\mathcal{F}$, we obtain a filtration
$0=\mathcal{U}_{n}\subseteq\cdots\subseteq\mathcal{U}_{0}=\mathcal{U}$
of coherent sheaves on $E$. Identifying the sub-quotients with their restrictions to $W$, for each $i$ one gets
${\mathcal{U}}_{i}/{\mathcal{U}}_{i+1}\cong\mathcal{U}|_{W}\otimes_{\mathcal{O}_{W}}\mathcal{L}^{i}$
where $\mathcal{L}=\mathcal{I}/{\mathcal{I}}^{2}$ is the line bundle that is the restriction of $\mathcal{I}$ to $W$. Observe that $\mathcal{L}$ is ample because $\mathcal{L}^{n}\cong{\mathcal{I}}^{n}/{\mathcal{I}}^{n+1}$ is the restriction of $\mathcal{O}_{B}(1)$ to $W$.
Set $\mathcal{G}\coloneqq({\mathcal{U}}_{0}/{\mathcal{U}}_{1})(-1)$, again viewed as a sheaf on $W$. Since $\dim W=1$ one has that $\dim\mathcal{G}=1$ as well and hence the function
$\chi(\mathcal{G}\otimes\mathcal{L}^{j})=\mathrm{rank}_{k}\operatorname{H}^{0}(W,\mathcal{G}\otimes\mathcal{L}^{j})-\mathrm{rank}_{k}\operatorname{H}^{1}(W,\mathcal{G}\otimes\mathcal{L}^{j})$
is strictly increasing in $j$. Indeed $\chi(\mathcal{G}\otimes\mathcal{L}^{j})$ is a polynomial function in $j$ of degree at most one, by  [19, Lemma 33.4.5.1], and since $\mathcal{L}$ is ample and $\dim\mathcal{G}=1$, this polynomial must be linear with positive slope. See [6, Propositions 23.78, 23.79] for more precise results.
The filtration of $\mathcal{U}$ above gives an injection and a surjection:
$\mathcal{G}\otimes\mathcal{L}^{n-1}\hookrightarrow{\mathcal{U}}(-1)\quad\text{and}\quad{\mathcal{U}}(-1)\twoheadrightarrow\mathcal{G}\,.$
Since $\mathcal{U}$ is an Ulrich sheaf, ${\mathcal{U}}(-1)$ has no cohomology; see Remark 2.1. Combined with the fact that $\dim W=1$ one gets
$\operatorname{H}^{0}(W,\mathcal{G}\otimes\mathcal{L}^{n-1})=0=\operatorname{H}^{1}(W,\mathcal{G})$
and therefore that
$\chi(\mathcal{G}\otimes\mathcal{L}^{n-1})\leq 0\leq\chi(\mathcal{G})\,.$
Since $n\geq 2$, this contradicts the fact that $\chi(\mathcal{G}\otimes\mathcal{L}^{j})$ is strictly increasing.
This completes the proof when $\dim R=2$.
It remains to explain the reduction to this case. So suppose $\dim R\geq 3$ and let $U$ be an Ulrich $R$-module. By enlarging the residue field if needed we can assume it is infinite. One can then find a system of parameters $\boldsymbol{x}\coloneqq x_{1},\dots,x_{d}$ of $R$ that is a regular sequence on $U$ and such that $\boldsymbol{x}U=\mathfrak{m}U$. Let
$R^{\prime}=R/(x_{1},\dots,x_{d-2})\quad\text{and}\quad U^{\prime}=U/(x_{1},\dots,x_{d-2})U.$
It is clear that the $R^{\prime}$-module $U^{\prime}$ is Ulrich. So it suffices to show the hypotheses still hold for the ring $R^{\prime}$.
Let $B$ and $B^{\prime}$ denote the blowups of $\operatorname{Spec}R$ and $\operatorname{Spec}R^{\prime}$ at their respective closed points, and let $E$ and $E^{\prime}$ be the corresponding exceptional fibers. It is straightforward to verify that the following diagram of rings is co-cartesian:
[…]
It thus follows that the square below is cartesian:
[…]
Our hypothesis is that the ideal $\mathcal{J}$ cutting $E$ out of $B$ satisfies $\mathcal{J}=\mathcal{I}^{n}$ for some locally principal sheaf of ideals $\mathcal{I}$ on $B$ and integer $n\geq 2$. Let $\mathcal{I}^{\prime}$ and $\mathcal{J}^{\prime}$ denote the pull-back of $\mathcal{I}$ and $\mathcal{J}$, respectively, to $B^{\prime}$. Since the square above is cartesian, $\mathcal{J}^{\prime}$ coincides with the ideal cutting out $E^{\prime}$ from $B^{\prime}$ and hence it is locally generated by a non-zero-divisor. Since $\mathcal{J}=\mathcal{I}^{n}$, we have $\mathcal{J}^{\prime}=(\mathcal{I}^{\prime})^{n}$, and since $\mathcal{I}$ is locally principal, so is $\mathcal{I}^{\prime}$. Thus, all the assumptions are valid for $R^{\prime}$, as desired. ∎
Here is one source of rings satisfying the hypotheses of the preceding result.
Section rings
The section ring of a line bundle $\mathcal{L}$ over a scheme $V$ is:
$\operatorname{S}(V,\mathcal{L})\coloneqq\bigoplus_{j\geqslant 0}\Gamma(V,\mathcal{L}^{j})\,.$
When $k$ is a field, $V$ is a proper $k$-scheme, and $\mathcal{L}$ is ample, $\operatorname{S}(V,\mathcal{L})$ is a finitely generated graded $k$-algebra.
Corollary 3.2.
Let $V$ be a projective $k$-scheme with $\dim V\geq 1$, let $\mathcal{L}$ be an ample line bundle on $V$, and $S=\operatorname{S}(V,\mathcal{L})$ the section ring of $\mathcal{L}$. Suppose there is an integer $a\geq 2$ so that the following conditions hold:
  1.  (1)
    $S_{0}=k$ and $S_{j}=0$ for $11\leq j\leq a-$ 1;
  2.  (2)
    the multiplication map $S_{a}\otimes_{k}S_{j}\to S_{a+j}$ is surjective for all $j\geq a$.
Then for $\mathfrak{m}=S_{\geqslant 1}$, neither the local ring $S_{\mathfrak{m}}$ nor its $\mathfrak{m}$-adic completion admits an Ulrich module.
Proof.
We claim that as schemes over $\operatorname{Spec}S$ one has
$\operatorname{Bl}_{\mathfrak{m}}(S)\cong\operatorname{\mathbf{Spec}}_{V}(\mathrm{Sym}(\mathcal{L}))\quad\text{where $\mathrm{Sym}(\mathcal{L})$ is the sheaf of algebras $\bigoplus_{i\geqslant 0}\mathcal{L}^{i}$.}$
To verify this, we first verify by an induction on $j$ that $\mathfrak{m}^{j}=S_{\geqslant ja}$ for all $j\geq 0$.
The base case $j=1$ is clear. Assuming the desired equality holds for some $j\geq 1$, one gets equalities
$\mathfrak{m}^{j+1}=\mathfrak{m}\mathfrak{m}^{j}=S_{\geqslant a}\cdot S_{\geqslant aj}=S_{\geqslant(j+1)a}$
where the second one is by the induction hypothesis, and the last one is immediate from condition (2) in the statement.
The equality $\mathfrak{m}^{j}=S_{\geqslant ja}$ justifies the first isomorphism below:
$\operatorname{Bl}_{\mathfrak{m}}(S)\cong\operatorname{Proj}\big{(}\bigoplus_{j\geqslant 0}S_{\geqslant ja}t^{j}\big{)}\cong\operatorname{Proj}\big{(}\bigoplus_{j\geqslant 0}S_{\geqslant j}t^{j}\big{)}\cong\operatorname{\mathbf{Spec}}_{V}(\mathrm{Sym}(\mathcal{L}))\,,$
The second one holds as $\oplus_{j\geqslant 0}S_{\geqslant ja}t^{j}$ is the $a$-th Veronese subring of $\oplus_{j\geqslant 0}S_{\geqslant j}t^{j}$, and the last isomorphism holds because $\mathcal{L}$ is ample; see [7, Section 8.7.3] or [11, Paragraph 6.2.1]. These isomorphisms are compatible with projections to $\operatorname{Spec}S$.
This justifies the claim about the blowup.
The exceptional fiber $E$ of the blowup $\operatorname{Bl}_{\mathfrak{m}}(S)\to\operatorname{Spec}S$ is the subscheme of $\operatorname{Bl}_{\mathfrak{m}}(S)$ cut out by $\mathcal{L}^{a}$, that is to say that
$E\cong\operatorname{\mathbf{Spec}}\left(\frac{\mathrm{Sym}(\mathcal{L})}{\mathcal{L}^{a}\,\mathrm{Sym}(\mathcal{L})}\right)\,.$
As $V$ is the subscheme of $\operatorname{Bl}_{\mathfrak{m}}(S)$ cut out by $\mathcal{L}$ and $a\geq 2$, we can apply Theorem 3.1 to get the desired result. ∎
Remark 3.3.
Conditions (1) and (2) imply that $S$ is generated as a $k$-algebra by its components in degrees $[a,2a-1]$. However the latter condition is weaker: the $k$-algebra $k[x,y]$ with $|x|=2$ and $|y|=3$ is generated by its components in degrees $[2,3]$, but the map $S_{2}\otimes_{k}S_{4}\to S_{6}$ is not surjective for $y^{2}$ is not in the image.
Corollary 3.2 brings up the question: When is a noetherian graded ring over a field the section ring of an ample line bundle on a projective scheme? Here is one answer; it is extrapolates from [11, Proposition 2.1], where the result is stated for normal domains. The argument does not require these conditions, and we sketch a proof, to convince ourselves.
Lemma 3.4.
Let $k$ be a field and $S$ a noetherian $\mathbb{N}$-graded $k$-algebra with $S_{0}=k$ and $0ptS\geq 2$. Then $S$ is the section ring of an ample line bundle on a projective scheme over $k$ if and only if there exist homogeneous elements $x_{0},\dots,x_{d}$ in $S$ such that $\sqrt{(x_{0},\dots,x_{d})}=S_{\geqslant 1}$ and the $\mathbb{Z}$-graded rings $S_{x_{i}}$ has a unit of degree $1$.
Proof.
Suppose $S=\operatorname{S}(X,\mathcal{L})$. Choose an integer $m\gg 0$ such that $\mathcal{L}^{m}$ and $\mathcal{L}^{m+1}$ are very ample (see [9, Theorem II, 7.6] and [14, Theorem 1.2.6]) and let $s_{1},\dots,s_{r}$ and $t_{1},\dots,t_{l}$ be bases of $\Gamma(X,\mathcal{L}^{m})$ and $\Gamma(X,\mathcal{L}^{m+1})$, respectively. One can check that the radical of the ideal generated by $x_{ij}=s_{i}t_{j}$, for $1\leq i\leq r$ and $1\leq j\leq l$, is $S_{\geqslant 1}$ and that $t_{j}^{2}/x_{ij}$ in $S_{x_{ij}}$ is a unit of degree 1.
For the converse, by computing global sections using Čech cover associated to the sequence $\boldsymbol{x}=(x_{0},\dots,x_{d})$, one gets an exact sequence
[…]
where $\mathfrak{m}\coloneqq S_{\geqslant 1}$, the homogeneous maximal ideal of $S$. Since $0ptS\geq 2$, we deduce that the map $\sigma$, which is a map of graded $k$-algebras, is an isomorphism. The hypothesis on $\boldsymbol{x}$ translates to the existence of an open cover
$\operatorname{Proj}S=D(x_{0})\cup\cdots\cup D(x_{d})$
such that one has an isomorphism of $\mathbb{Z}$-graded rings $S_{x_{i}}\cong[S_{x_{i}}]_{0}[t^{\pm 1}]$, where $|t|=1$. Using this it is easy to check that the natural map
$\widetilde{S(i)}\otimes\widetilde{S(j)}\longrightarrow\widetilde{S(i+j)}$
is an isomorphism for all integers $i,j$. Hence $\mathcal{L}\coloneqq\widetilde{S(1)}$ is invertible and $\widetilde{S(j)}\cong\mathcal{L}^{j}$. It follows that $S$ is the section ring of the ample line bundle $\mathcal{L}$ on $\operatorname{Proj}S$. ∎
Here is another special case of Theorem 3.1.
Corollary 3.5.
Let $(R,\mathfrak{m},k)$ be a noetherian local ring such that $\dim R\geq 2$. Let $\operatorname{Bl}_{\mathfrak{m}}(R)\to\operatorname{Spec}R$ be the blowup of $\{\mathfrak{m}\}$ and $E$ the exceptional fiber of the blowup. If $\operatorname{Bl}_{\mathfrak{m}}(R)$ is locally factorial (for instance, regular) and $E$ is irreducible but not reduced, then $R$ has no Ulrich modules.
Proof.
Set $W\coloneqq E_{\mathrm{red}}$. The hypothesis implies that it is an integral scheme and hence an integral Weil divisor on $\operatorname{Bl}_{\mathfrak{m}}(R)$. Since $\operatorname{Bl}_{\mathfrak{m}}(R)$ is locally factorial, we know that $W$ is a Cartier divisor and $[E]=n[W]$ for some $n\geq 2$, since $E$ is not reduced. Thus Theorem 3.1 applies. ∎
Our interest in the formulation above is that it applies also when $R$ is not itself a section ring. To illustrate this, we revisit the example constructed by Yhee [21].
Yhee’s example revisited
Let $k$ be a field, $S=k[\boldsymbol{x}]$ the polynomial ring over $k$ in indeterminates $\boldsymbol{x}=x_{1},\dots,x_{d}$, and set $\mathfrak{m}=(\boldsymbol{x})$. Fix an integer $n\geq 2$, set $J=\mathfrak{m}^{n}$ and $R=k\oplus J$, viewed as a subring of $S$. Then $R$ cannot be a section ring, for it is not $S_{2}$.
We claim that the ring $R$ satisfies the hypothesis of Corollary 3.5.
Indeed, the cokernel of the inclusion $\mathscr{R}_{J}(R)\subseteq\mathscr{R}_{J}(S)$ is $S/R$, and in particular finite dimensional as a $k$-vector space. Thus one gets the first isomorphism below:
$\operatorname{Bl}_{J}(R)\xleftarrow{\ \cong\ }\operatorname{Bl}_{J}(S)\cong\operatorname{Bl}_{\mathfrak{m}}(S)\cong\mathbb{P}^{d-1}_{S}\,.$
The second isomorphism holds as $\mathscr{R}_{J}(S)$ is the $n$th Veronese subring of $\mathscr{R}_{\mathfrak{m}}(S)$; this holds because $J=\mathfrak{m}^{n}$. In particular $\operatorname{Bl}_{J}(R)$ is regular. The exceptional fiber of the blowup is $E=\operatorname{Proj}(\operatorname{gr}_{J}(R))$ where
$\operatorname{gr}_{J}(R)=k\oplus\frac{\mathfrak{m}^{n}}{\mathfrak{m}^{2n}}\oplus\frac{\mathfrak{m}^{2n}}{\mathfrak{m}^{3n}}\oplus\cdots.$
Observe that the ideal sheaf, say $\mathcal{I}$, corresponding to $\mathfrak{m}^{n+1}\operatorname{gr}_{J}(R)$ is nilpotent, of order $n$, and nonzero; the latter holds because $n\geq 2$. So $E$ is not reduced. It is irreducible because $E/\mathcal{I}$ is the projective scheme associated to the $k$-algebra
$k\oplus\bigoplus_{i\geqslant 1}\frac{\mathfrak{m}^{ni}}{\mathfrak{m}^{ni+1}},$
which is isomorphic to the $n$th Veronese subring of $S$, so $E/\mathcal{I}\cong\mathbb{P}^{d-1}_{k}$.
4. Examples: Complete intersections
In this section we construct complete intersection local rings, some of which are domains, that have no Ulrich modules. Let $k$ be a field and consider the $k$-algebra
$S\coloneqq\frac{k[s,t,x,y,z]}{(s^{2}-x^{3},st-f(x,y,z),t^{2}-z^{3})}$
where $f(x,y,z)$ is a homogeneous polynomial in $k[x,y,z]$ of the form
$f(x,y,z)=y^{3}+f_{2}(x,z)y^{2}+f_{1}(x,z)y+f_{0}(x,z)$
where each $f_{i}(x,z)$ is in $k[x,z]$. Set the degree of $x,y,z$ to $2$ and that of $s,t$ to $3$. Then $S$ is an $\mathbb{N}$-graded ring, with homogeneous maximal ideal $\mathfrak{m}\coloneqq S_{\geqslant 2}$.
Theorem 4.1.
For any choice of $f(x,y,z)$ satisfying the condition above neither the local ring $S_{\mathfrak{m}}$ nor its $\mathfrak{m}$-adic completion has an Ulrich module. They are both complete intersections of dimension $2$.
Proof.
It suffices to verify that the graded $k$-algebra $S$ has the following properties:
  1.  (a)
    It is a complete intersection ring, with $\dim S=2$;
  2.  (b)
    It satisfies the conditions (1) and (2) of Corollary 3.2, with $a=2$;
  3.  (c)
    It is the section ring of an ample line bundle on a projective curve.
Then Corollary 3.2 yields the desired conclusion.
Proof of (a): It is clear that $\operatorname{gr}_{\mathfrak{m}}(S)$ is a quotient of the ring
$\frac{k[s,t,x,y,z]}{(s^{2},st,t^{2},x^{3}z^{3}-f(x,y,z)^{2})}\,.$
Since $f(x,y,z)$ is monic in $y$ we get that $\dim S\leq 2$. On the other hand, $S$ is defined by three relations, so it must be a complete intersection ring, with $\dim S=2$.
Proof of (b): Evidently $S_{0}=k$ and $S_{1}=0$. Let $R$ be the second Veronese subring of $S$. It is clear from the relations defining $S$ that $R$ is the $k$-sub-algebra of $S$ generated by $x,y,z$. As an $R$-module, $S$ is generated by $1,s,t$. It follows that $S_{2}\cdot S_{j}=S_{j+2}$ for $j\geq 2$.
Proof of (c): We use the criterion from Lemma 3.4. The elements $x,z$ form a homogeneous system of parameters for $S$. The element $s/x$ is of degree $1$ in $S_{x}$, and invertible, with inverse $s/x^{2}$. By symmetry, $t/z$ is a unit of degree $1$ in $S_{z}$.
This completes the proof of the theorem. ∎
The preceding theorem provides a family of complete intersections with no Ulrich modules. Special values of $f(x,y,z)$ yield rings with further interesting properties.
The case $f(x,y,z)=y^{3}+x^{2}z$
Then the ring in question is a domain
$S\coloneqq\frac{k[s,t,x,y,z]}{(s^{2}-x^{3},st-(y^{3}+x^{2}z),t^{2}-z^{3})}.$
For $k=\mathbb{Q}$, this claim can be verified using, for instance, Macaulay2. Here is a sketch of an argument for general $k$: Since $x$ is a parameter in $S$, it is not a zero-divisor, so it suffices to check that the ring $S[x^{-1}]$ is a domain. One has
$S[x^{-1}]=\frac{A[x^{-1},s]}{(s^{2}-x^{3})}\quad\text{where}\quad A=\frac{k[x,y,z]}{((y^{3}+x^{2}z)^{2}-x^{3}z^{3})}$
It thus suffices to verify that $A$ is a domain and $x^{3}$ does not have a square-root in the fraction field of $A$.
The property that $A$ is a domain translates to the irreducibility of the polynomial $(y^{3}-x^{2}z)^{2}-x^{3}z^{3}$. Specializing $y=1$, it suffices to verify that the polynomial
$x^{4}z^{2}-x^{3}z^{3}-2x^{2}z+1$
is irreducible. This can be done easily using the Newton polytope associated to this polynomial, as described in Gao [5]: The polytope is a triangle with vertices
$v_{0}=(0,0),v_{1}=(3,3),v_{2}=(4,2).$
The greatest common divisor of the coordinates of $v_{0}-v_{1}$ and $v_{0}-v_{2}$ is $1$, so the polytope is integrally indecomposable, by [5, Corollary 4.5]. We conclude that polynomial in question is irreducible; see [5, p. 507].
Finally, if $x^{3}$ has a square-root in the fraction field of $A$, then so is $x$. Thus there exist nonzero elements $a,b\in A$ such that $a^{2}x=b^{2}$. Since $A$ can be standard graded, with degrees of $x,y,z$ equal to $1$, this is not possible since the highest degree term in $a^{2}x$ has odd degree while the highest degree term in $b^{2}$ has even degree.
The case $f(x,y,z)=y^{3}$ and $k$ is algebraically closed
The complete intersection in question is then
$S\coloneqq\frac{k[s,t,x,y,z]}{(s^{2}-x^{3},st-y^{3},t^{2}-z^{3})}$
and this is interesting, for it is closely connected to Yhee’s construction [21] of domains (not Cohen-Macaulay) that have no Ulrich modules.
Indeed, according to Macaulay2, the ring $S$ has three minimal prime ideals. Evidently, for any of minimal prime $\mathfrak{p}$ the domain $S/\mathfrak{p}$ (or rather it localization at the maximal ideal) cannot have Ulrich modules, for any Ulrich module over it would be Ulrich over $S_{\mathfrak{m}}$. Again according to Macaulay2, one of the prime ideals is
$\mathfrak{p}=(s^{2}-x^{3},st-xyz,z^{3}-t^{2},y^{2}-xz,sz^{2}-xyt,yzs-x^{2}t)$
as an ideal in the underlying polynomial ring. The assignment
$s\mapsto u^{3},t\mapsto v^{3},x\mapsto u^{2},y\mapsto uv,z\mapsto v^{2}$
induces an isomorphism of $k$-algebras
$S/\mathfrak{p}\cong k[u^{2},uv,v^{2},u^{3},v^{3}]\subset k[u,v].$
This subring on the right differ from the subring $k+(u,v)^{2}$ considered in [21], as discussed in the previous section, only in containing $u^{2}v,uv^{2}$, and one can argue as in [21] to deduce that $S/\mathfrak{p}$ does not have Ulrich modules.
The same analysis carries over to the quotient of $S$ by the other minimal primes. Summing up, the complete intersection ring $S$ above can be thought of being obtained by gluing three copies of the (non Cohen-Macaulay) ring constructed by Yhee in [21], as discussed in the previous section.
Remark 4.2.
More generally, consider the $k$-algebra
$S\coloneqq\frac{k[s_{1},\dots,s_{n},x_{ij}\mid 1\leq i\leq j\leq n]}{(s_{i}s_{j}-x_{ij}^{3}\mid 1\leq i\leq j\leq n)}\,.$
Then $S_{\mathfrak{m}}$ is a complete intersection of dimension $n$, and an argument similar to the proof of Theorem 4.1 shows that $S_{\mathfrak{m}}$ does not admit Ulrich modules. Moreover the ideal $(x_{11},\dots,x_{nn})$ is a minimal reduction of $\mathfrak{m}$; it is easy to check the other variables are integral over this ideal. Thus
$e(S_{\mathfrak{m}})$ $=\operatorname{length}\left(\frac{S}{(x_{11},\dots,x_{nn})}\right)$
$=\operatorname{length}\left(\frac{k[s_{1},\dots,s_{n},x_{ij}\mid 1\leq i<j\leq n]}{(s_{1}^{2},\dots,s_{n}^{2},s_{i}s_{j}-x_{ij}^{3}\mid 1\leq i<j\leq n)}\right)$
$=2^{n}\cdot 3^{\binom{n}{2}}.$
where the last equality follows because the ring in question is a finite free extension of $k[s_{1},\dots,s_{n}]/(s_{1}^{2},\dots,s_{n}^{2})$ of degree $3^{\binom{n}{2}}$. Note that $S_{\mathfrak{m}}$ is far from a strict complete intersection, as the product of the $\mathfrak{m}$-adic orders of the defining equations is $2^{\binom{n+1}{2}}$.
One thus gets a family of complete intersections with no Ulrich modules, and for which the difference between the orderand its multiplicity is arbitrarily large.
5. Examples: Gorenstein normal domains
In this section we construct Gorenstein normal domains with no Ulrich modules.
Let $C$ be a smooth non-hyperelliptic curve of genus $g\geq 3$ (e.g. a smooth plane quartic), and let $\mathcal{L}$ be a non-effective theta characteristic on $C$. In other words,
$\mathcal{L}^{\otimes 2}\cong\omega_{C}\quad\text{and}\quad\Gamma^{0}(C,\mathcal{L})=0\,.$
It is a classical result [16] (see also [1]) that non-effective theta characteristic exists on any smooth curve. For example, one can take
$C=\operatorname{Proj}\frac{\mathbb{C}[x,y,z]}{(x^{4}+y^{4}+z^{4})}\,.$
There is an ample line bundle $\mathcal{L}$ on $C$ such that $\mathcal{L}^{2}=\omega_{C}$ and $\Gamma(C,\mathcal{L})=0$. In fact, for any plane quartic, there are $642^{2g}=$ 64 theta characteristics, but only $28$ of them are effective (corresponding to the $28$ bitangents); see [9, Exercise IV, 2.3(h)].
Theorem 5.1.
Set $S\coloneqq\operatorname{S}(C,\mathcal{L})$ and let $\mathfrak{m}$ be the homogeneous maximal ideal of $S$. Then neither the local ring $S_{\mathfrak{m}}$ nor its completion at $\mathfrak{m}$ admits an Ulrich module. They are both Gorenstein and normal.
Proof.
To begin with $S$ is normal because it is the section ring of smooth projective variety, with respect to an ample invertible sheaf; see [18, Section 1]. It is also Gorenstein by [13, Proposition 3.14(4)]. These conditions are then inherited by $S_{\mathfrak{m}}$ and its completion. For the rest, we show that the conditions in Corollary 3.2 are satisfied with $a=2$.
Indeed, since $C$ is not hyperelliptic, the canonical ring $\operatorname{S}(\omega_{C})$ is generated in degree 1 by a classical theorem of Max Noether. As $\mathcal{L}^{\otimes 2}\cong\omega_{C}$, we deduce that the multiplication map $\mathrm{Sym}^{m}(S_{2})\to S_{2m}$ is surjective for all integer $m$, hence $S_{2}\otimes_{k}S_{j}\to S_{2+j}$ is also surjective for all even integer $j\geq 2$. On the other hand, the line bundle $\mathcal{L}^{\otimes 3}$ on $C$ is Castelnuovo-Mumford 00-regular (with respect to the very ample line bundle $\omega_{C}$). To see this, it suffices to note that $\chi(\mathcal{L})=0$ by Riemann-Roch, and hence
$\operatorname{H}^{1}(C,\mathcal{L}^{\otimes 3}\otimes\omega_{C}^{-1})=\operatorname{H}^{1}(C,\mathcal{L})=\operatorname{H}^{0}(C,\mathcal{L})=0\,.$
By [14, Theorem 1.8.3], it follows that the multiplication map $S_{3}\otimes\mathrm{Sym}^{m}(S_{2})\to S_{2m+3}$ is also surjective for all integer $m$, which in turn implies the surjectivity of $S_{2}\otimes_{k}S_{j}\to S_{2+j}$ for all odd integer $j\geq 3$. This verifies the condition (2) from Corollary 3.2, while condition (1) is clear from the choice of $\mathcal{L}$. ∎
Remark 5.2.
With $g$ the genus of the curve $C$, a direct computation gives that the Hilbert series of the ring $S$ in Theorem 5.1 is
$\frac{(1-2t+(g+1)t^{2}-2t^{3}+t^{4})}{(1-t)^{2}}$
If $S$ were a complete intersection, the numerator would be product of cyclotomic polynomials. One can check that this is not the case.
Remark 5.3.
The ring $S_{\mathfrak{m}}$ in Theorem 5.1 is not a rational singularity. In fact, rational surface singularities have minimal multiplicities (that is to say, the multiplicity is one less than the embedding dimension, see [17, Theorem on page 94]), and such rings admit Ulrich modules, see [2, Proposition 2.5]. It is an interesting question to find an example of a rational singularity with no Ulrich modules.
Acknowledgements.
Many thanks to Kevin Tucker for helpful conversations concerning this work. The authors were partly supported by National Science Foundation (NSF) grants DMS-2001368 (SBI); DMS-2302430 (LM); DMS-1952366 (LM); DMS-2200732 (MW); DMS-2240926 (ZZ), and DMS-2234736 (ZZ). Ma was also partly supported by a fellowship from the Sloan Foundation and by a grant from the Institute for Advanced Study School of Mathematics. Zhuang was also partly supported by a Clay research fellowship, as well as a Sloan fellowship.
This material is based upon work supported by the NSF under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while the first three authors were in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Spring 2024 semester.